### v2

@@ 3 Terminal conditions with bounded Malliavin derivative
This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.6) under the assumption that the driver $g$ is Lipschitz continuous and the terminal condition $\xi$ is Malliavin differentiable. These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques. Therefore, we initially work under stronger regularity assumptions to facilitate this reduction. For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition and condition on $\xi$ will be introduced in Propisiton 3.1 below.
  • $({\rm H_{lip}})$
@@
The existence and uniqueness of a solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ to the BSDE (2.6) under assumptions $(\rm H_{0})$$({\rm H_{lip}})$ is a well-established result, as shown in Theorem 3.1 of Buckdahn2009MFBSDE. In the following proposition, we demonstrate that the process $Y$, which lies in $\mathbb{S}^{\infty}$, is bounded by a constant that is independent of the Lipschitz constant $L_{z}$. More remarkably, we show that the same independence holds for the process $Z$, provided the terminal condition $\xi$ admits uniformly bounded Malliavin derivatives.
@@
If $(\rm H_{0})$} and $({\rm H_{lip}})$ hold, then the unique solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ of the Lipschitz mean-field BSDE satisfies the following properties.
@@
    $\left\|Z\right\|_{\mathbb{S}^{\infty}}\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}},$ (3.8)
@@
Unlike standard Lipschitz BSDE, the comparison theorem cannot be applied here (see Theorem 3.2 in Buckdahn2009MFBSDE). Observe that (3.7) can be derived following the line of proof of Proposition 4.4. In this Lipschitz context, the proof is simpler since no BMO martingale arguments are necessary.
@@
Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$. Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]$ is measurable to $mathcal{F_{s}}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
@@
Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.8. ∎
@@ 4 A priori estimates and stability for BSDEs
In this section, we will work under the assumptions $(\rm H_{0})$} and $(\rm H_{q})$} to establish the stability of the mean-field BSDE (2.6), as stated in Theorem 4.6. The strategy of the proof is as follows. We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.6). These estimates, together with fundamental properties of BMO martingales, lead to the stability result. We begin with a useful estimate that will play an important role in the analysis.
@@
Under assumptions $(\rm H_{0})$} and $(\rm H_{q})$}, we have the following estimate:
@@
[…]
for any $(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}$. Taking expectations with respect to $\tilde{\mathbb{P}}$ yields:
$\tilde{\mathbb{E}}\left|g(t,Z_{t},\tilde{Z}_{t})\right|\leq K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}+\frac{3K_{z}}{2}\left|Z_{t}\right|^{2}+\frac{3K_{\tilde{z}}}{2}\tilde{\mathbb{E}}\left|\tilde{Z}_{t}\right|^{2}.$ (4.9)
@@
Suppose that (H0) and (Hq) hold. Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a solution to the mean-field BSDE (2.6). Then the stochastic process $\left(\int_{0}^{t}Z_{s}\!\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale. Moreover, its BMO norm satisfies the following bound:
$\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}^{2}\leq(1\vee T)\big(4K_{0}+K_{z}+K_{\tilde{z}}\big)\,\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\left(\frac{K_{\tilde{z}}}{K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\right).$ (4.10)
@@
The argument follows from arguments in Proposition 2.1 of briand2008stobsdes with modified $C^{2}$ function
$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad c\in\mathbb{R},x\in\mathbb{R}.$
In Step 1 we express the $\mathrm{BMO}$ norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
@@
$0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{3K_{z}}\,e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.11)
Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.9 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
@@
$\leq\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(\frac{3K_{z}}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
$\leq\varphi(\xi)-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
where $K:=K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}$. Taking $\mathbb{E}[\cdot\mid\mathcal{F}_{\tau}]$ and using (4.11) together with $|\varphi(x)|\leq C_{Y}|x|$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded on the range of $Y$), we get
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
$\leq C_{Y}\,\mathbb{E}\!\big[|\xi|\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}KT+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds.$
we thus obtain
$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.12)
@@
Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality, assume $\alpha$ satisfies
$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-2}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
by enlarging $K_{z}$ or $K_{\tilde{z}}$. Define positive constants
$c_{1}:=\frac{K_{\tilde{z}}}{2K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+\frac{1}{2},\qquad c_{2}:=\tfrac{1}{2}e^{3K_{z}}.$
Note that, by definition of $\alpha$, $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)$. By the choice of $\alpha$, it is easy to verify $c_{1}<c_{2}$. Pick any $c\in(c_{1},c_{2})$ and consider the same $\varphi$ at in Step 1 with a different $c$ value, so that $\varphi(0)=0$ and, for all $x$,
$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}\geq 0.$
We also have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 with $\tau=0$ gives
$\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}ds+\frac{3K_{\tilde{z}}}{2}\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,\mathbb{E}|Z_{s}|^{2}ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
Taking expectations and using $|\varphi^{\prime}(x)|=\frac{e^{3K_{z}x}-2c}{3K_{z}}\leq\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}=C_{Y}$ for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$,
$\varphi(Y_{0})\leq\mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T-c\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]+\frac{3K_{\tilde{z}}}{2}\,C_{Y}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right].$
By the definition of $c_{1}$ and the fact that $c>c_{1}=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}<c-\tfrac{1}{2}$. Hence
$\frac{1}{2}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]\ \leq\ \mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T\ \leq\ \frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\,\big(\left\|\xi\right\|_{L^{\infty}}+KT\big).$ (4.13)
@@
Finally, plugging (4.13) into (4.12) and taking the essential supremum over $\tau$ yields an explicit $\mathrm{BMO}$ bound for $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ in terms of $\|Y\|_{\mathbb{S}^{\infty}}$ and the model parameters. This completes the proof. ∎
@@
Next, suppose the solution $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$. Then, for any solution pair $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the quadratic mean-field BSDE, we derive the following proposition showing that $Y\in\mathbb{S}^{\infty}$ which is independent of the BMO norm $\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}$. We introduce a simple but useful lemma which will be used later on when we work under equivalent probability measures.
Lemma 4.3.
Let $\mathbb{P}$ and $\mathbb{Q}$ be equivalent probability measure on $(\Omega,\mathcal{F})$. For any random variable $X$,
$\|X\|_{L^{\infty}(\mathbb{P})}=\|X\|_{L^{\infty}(\mathbb{Q})}.$
In particular, $X$ is bounded $\mathbb{P}$-a.s. iff it is bounded $\mathbb{Q}$-a.s., with the same bound.
Proof.
Set $K:=\|X\|_{L^{\infty}(\mathbb{P})}$. For every $\varepsilon>0$, $\mathbb{P}(|X|>K+\varepsilon)=0$. Since $\mathbb{Q}\ll\mathbb{P}$, also $\mathbb{Q}(|X|>K+\varepsilon)=0$, so $\|X\|_{L^{\infty}(\mathbb{Q})}\leq K$. By symmetry ($\mathbb{P}\ll\mathbb{Q}$), the reverse inequality holds. ∎
@@
Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
@@
The proof relies on a linearization argument combined with the BMO property. Using Remark 2.6, without the loss of generality, we assume
@@
Then, the BSDE (2.6) can be rewritten as

### v3

@@ 3 Terminal conditions with bounded Malliavin derivative
This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.6) under the assumption that the driver $g$ is Lipschitz continuous and the terminal condition $\xi$ is Malliavin differentiable. These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver of quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques. Therefore, we initially work under stronger regularity assumptions to facilitate this reduction. For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition and condition on $\xi$ will be introduced in Propisiton 3.1 below.
  • $(\mathrm{H}_{\mathrm{lip}})$
@@
The existence and uniqueness of a solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ to the BSDE (2.6) under assumptions $(\mathrm{H}_{0})$$(\mathrm{H}_{\mathrm{lip}})$ is a well-established result, as shown in Theorem 3.1 of Buckdahn et al. (2009b). In the following proposition, we demonstrate that the process $Y$, which lies in $\mathbb{S}^{\infty}$, is bounded by a constant that is independent of the Lipschitz constant $L_{z}$. More remarkably, we show that the same independence holds for the process $Z$, provided the terminal condition $\xi$ admits uniformly bounded Malliavin derivatives.
@@
If $(\mathrm{H}_{0})$ and $(\mathrm{H}_{\mathrm{lip}})$ hold, then the unique solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ of the Lipschitz mean-field BSDE satisfies the following properties.
@@
    $\left\|Z\right\|_{\mathbb{S}^{\infty}}\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}}.$ (3.8)
@@
Unlike standard Lipschitz BSDE, the comparison theorem cannot be applied here (see Theorem 3.2 in Buckdahn et al. (2009b)). Observe that (3.7) can be derived following the line of proof of Proposition 4.3. In this Lipschitz context, the proof is simpler since no BMO martingale arguments are necessary.
@@
Under assumption $(\mathrm{H}_{\mathrm{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$. Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]\in\mathcal{F}_{s}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new measure $\mathbb{Q}$:
@@
Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish (3.8). ∎
@@ 4 A priori estimates and stability for BSDEs
In this section, we establish stability estimates for the mean-field BSDE (2.6) under two different quadratic structures of the generator. More precisely, we prove Theorem 4.6 under the unified quadratic condition $(\mathrm{H}_{q}^{\mathrm{unif}})$ (together with the additional terminal condition assumption $(\mathrm{H}_{\xi})$), and Theorem 4.7 under the separable quadratic condition $(\mathrm{H}_{q}^{\mathrm{sep}})$. Both results yield the same stability estimate in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. The overall strategy in both regimes is as follows. We first derive a priori estimates for any solution $(Y,Z)$ to (2.6): an $\mathbb{S}^{\infty}$ bound for $Y$ and a BMO bound for $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$. These estimates, together with fundamental properties of BMO martingales, lead to the stability results. We begin with a growth bound for the generator, which will be used repeatedly throughout the section.
@@
Under assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$}, we have the following estimate:
@@
for any $(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}$, where $K_{0}:=\sup_{t\in[0,T]}|g(t,0,0)|$. Taking expectations with respect to $\tilde{\mathbb{P}}$ yields:
$\tilde{\mathbb{E}}\,|g(t,Z_{t},\tilde{Z}_{t})|\leq K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}+\Big(2K_{z}+\frac{1}{2}K_{\tilde{z}}\Big)|Z_{t}|^{2}+\Big(2K_{\tilde{z}}+\frac{1}{2}K_{z}\Big)\tilde{\mathbb{E}}|\tilde{Z}_{t}|^{2},$ (4.9)
where we used symmetry $\tilde{\mathbb{E}}|\tilde{Z}_{t}|={\mathbb{E}}|Z_{t}|$, together with $|Z_{t}|\,{\mathbb{E}}|Z_{t}|\leq\frac{1}{2}|Z_{t}|^{2}+\frac{1}{2}{\mathbb{E}}|Z_{t}|^{2}$.
@@
Suppose that $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold. Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a solution to the mean-field BSDE (2.6). Then the stochastic process $\big(\int_{0}^{t}Z_{s}\cdot dW_{s}\big)_{0\leq t\leq T}$ defines a BMO martingale. Moreover, its BMO norm satisfies the bound
$\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{\mathrm{BMO}}^{2}\leq(1\vee T)\,(4K_{0}+K_{z}+K_{\tilde{z}})\,\frac{e^{(4K_{z}+K_{\tilde{z}})(1+\|Y\|_{\mathbb{S}^{\infty}})}}{4K_{z}+K_{\tilde{z}}}\left(\frac{K_{z}+4K_{\tilde{z}}}{4K_{z}+K_{\tilde{z}}}\,e^{(4K_{z}+K_{\tilde{z}})\|Y\|_{\mathbb{S}^{\infty}}}+1\right).$ (4.10)
@@
The argument follows Proposition 2.1 of Briand and Confortola (2008), with constants adapted to the growth estimate (4.9). Define
[…]
Recall that, under $(\mathrm{H}_{0})$} and $(\mathrm{H}_{q}^{\mathrm{unif}})$}, the estimate (4.9) reads
$\widetilde{\mathbb{E}}\,|g(t,Z_{t},\tilde{Z}_{t})|\leq K+A|Z_{t}|^{2}+B\,\widetilde{\mathbb{E}}|\tilde{Z}_{t}|^{2},\qquad t\in[0,T].$
Define the $C^{2}$ function with parameter $c\in\mathbb{R}$
$\varphi(x):=\frac{e^{\lambda|x|}-1-2c\lambda|x|}{\lambda^{2}},\qquad x\in\mathbb{R},$
where $\lambda:=2A=4K_{z}+K_{\tilde{z}}$. In Step 1 we express the BMO norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$. Then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{\lambda|x|}-1}{\lambda}\,\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{\lambda|x|},\qquad\varphi^{\prime\prime}(x)-\lambda\,|\varphi^{\prime}(x)|\equiv 1.$
@@
$0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{\lambda}\,e^{\lambda\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.11)
Let $\tau$ be any stopping time. Applying Itô’s formula to $\varphi(Y)$ on $\tau,T]$ and using (4.9) together with $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
@@
$\leq\varphi(\xi)+\int_{\tau}^{T}\Big(A|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}\,ds+\int_{\tau}^{T}|\varphi^{\prime}(Y_{s})|\Big(K+B\,\mathbb{E}|Z_{s}|^{2}\Big)\,ds$
$\hskip 56.9055pt-\int_{\tau}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
Since $\lambda=2A$ and $\varphi^{\prime\prime}-\lambda|\varphi^{\prime}|\equiv 1$, we have
$A|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})=-\tfrac{1}{2}\big(\varphi^{\prime\prime}(Y_{s})-2A|\varphi^{\prime}(Y_{s})|\big)=-\tfrac{1}{2}.$
Taking $\mathbb{E}[\cdot\,|\,\mathcal{F}_{\tau}]$, using (4.11), and also $|\varphi(x)|\leq C_{Y}|x|$ on the range of $Y$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded there), we obtain
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\int_{\tau}^{T}|Z_{s}|^{2}\,ds\ \Big|\ \mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+BC_{Y}\int_{0}^{T}\mathbb{E}|Z_{s}|^{2}\,ds$
$\leq C_{Y}\,\|\xi\|_{L^{\infty}}+C_{Y}KT+BC_{Y}\int_{0}^{T}\mathbb{E}|Z_{s}|^{2}\,ds.$
Hence,
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\int_{\tau}^{T}|Z_{s}|^{2}\,ds\ \Big|\ \mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+B\int_{0}^{T}\mathbb{E}|Z_{s}|^{2}\,ds\right).$ (4.12)
@@
Step 2 (An explicit $\mathbb{H}^{2}$ bound). We assume (without loss of generality, by enlarging $K_{z}$ relative to $K_{\tilde{z}}$) that
$0<\alpha<\min\Big\{\tfrac{1}{2},\ \frac{e^{\lambda}-2}{e^{\lambda\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$ (4.13)
Define
$\widehat{C}_{Y}:=\frac{e^{\lambda(1+\|Y\|_{\mathbb{S}^{\infty}})}}{\lambda},\qquad c_{1}:=B\,\widehat{C}_{Y}+\frac{1}{2},\qquad c_{2}:=\frac{1}{2}e^{\lambda}.$
Then (4.13) implies $c_{1}<c_{2}$. Fix any $c\in(c_{1},c_{2})$ and consider the shifted $C^{2}$ function
$\varphi(x):=\frac{e^{\lambda(1+|x|)}-1-2c\lambda(1+|x|)}{\lambda^{2}},\qquad x\in\mathbb{R}.$ (4.14)
For $x\neq 0$ we have
$\varphi^{\prime}(x)=\frac{e^{\lambda(1+|x|)}-2c}{\lambda}\,\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{\lambda(1+|x|)},\qquad\text{and}\qquad\varphi^{\prime\prime}-\lambda|\varphi^{\prime}|\equiv 2c,$
and the same identities hold in the generalized Itô–Tanaka sense at $x=0$. Moreover, since $1+|x|\geq 1$, we have for all $x\in\mathbb{R}$,
$|\varphi^{\prime}(x)|=\frac{e^{\lambda(1+|x|)}-2c}{\lambda}>\frac{e^{\lambda(1+|x|)}-2c_{2}}{\lambda}\geq 0,$
and for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$,
$|\varphi^{\prime}(x)|\leq\frac{e^{\lambda(1+\|Y\|_{\mathbb{S}^{\infty}})}}{\lambda}=\widehat{C}_{Y}.$
Applying Itô’s formula to $\varphi(Y)$ on $[0,T]$ (as in Step 1, with $\tau=0$) and using (4.9) yields
$\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}\,ds+B\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,{\mathbb{E}}|Z_{s}|^{2}\,ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
Taking expectations and using $|\varphi^{\prime}(Y_{s})|\leq\widehat{C}_{Y}$ gives
$\varphi(Y_{0})\leq{\mathbb{E}}[\varphi(\xi)]+K\widehat{C}_{Y}T-c\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right]+B\widehat{C}_{Y}\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right].$
Since $c>c_{1}=B\widehat{C}_{Y}+\tfrac{1}{2}$, we can absorb the last term and obtain
$\frac{1}{2}\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right]\leq{\mathbb{E}}[\varphi(\xi)]+K\widehat{C}_{Y}T.$
Finally, since $\varphi(0)=0$ and $|\varphi^{\prime}(x)|\leq\widehat{C}_{Y}$ for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$, we have $|\varphi(x)|\leq\widehat{C}_{Y}|x|$, hence
$\frac{1}{2}\,{\mathbb{E}}\!\left[\int_{0}^{T}|Z_{s}|^{2}\,ds\right]\leq\widehat{C}_{Y}\big(\|\xi\|_{L^{\infty}}+KT\big).$ (4.15)
@@
Finally, plugging (4.15) into (4.12) and taking the essential supremum over $\tau$ completes the proof. ∎
@@
Next, suppose that the solution satisfies $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$. Then, for any solution pair $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the quadratic mean-field BSDE, we derive the following proposition showing that $Y\in\mathbb{S}^{\infty}$, with a bound that is independent of the BMO norm $\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}$.
@@
Suppose Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\mathrm{unif}})$} hold. Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and
@@
The proof relies on a linearization argument combined with the BMO property. Using Remark 2.9, without loss of generality we assume
@@
Then BSDE (2.6) can be rewritten as